A sample of 100 recorded deaths in the United States during the past year showed an average life span of 71.8 years. Assuming a population standard deviation of 8.9 years,

does this seem to indicate that the average life span today is greater than 70 years? Use a level of significance:

a. 1%

b. 5%

c. 10%

 


1
Expert's answer
2022-05-09T18:25:04-0400

a.

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 70"

"H_a:\mu>70"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a right-tailed test is "z_c = 2.3263."

The rejection region for this right-tailed test is "R = \{z: z > 2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{71.8-70}{8.9/\sqrt{100}}=2.0225"

Since it is observed that "z = 2.0225<2.3263=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(Z>2.0225)=0.021562," and since "p=0.021562>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is greater than 70, at the "\alpha = 0.01" significance level.


b.

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 70"

"H_a:\mu>70"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{71.8-70}{8.9/\sqrt{100}}=2.0225"

Since it is observed that "z = 2.0225>1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>2.0225)=0.021562," and since "p=0.021562<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 70, at the "\alpha = 0.05" significance level.


c.

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 70"

"H_a:\mu>70"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.10," and the critical value for a right-tailed test is "z_c = 1.2816."

The rejection region for this right-tailed test is "R = \{z: z > 1.2816\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{71.8-70}{8.9/\sqrt{100}}=2.0225"

Since it is observed that "z = 2.0225>1.2816=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>2.0225)=0.021562," and since "p=0.021562<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 70, at the "\alpha = 0.10" significance level.